OpenStudy (anonymous):

Differentiate the identity sin 2x = 2sin x cos x to develop the identity for cos 2x , in terms of sin x and cos x.

OpenStudy (anonymous):

like what are they trying to ask? i dont get it

OpenStudy (anonymous):

take the derivatives of sin(2x) and 2sin x cos x, then set the two equal to each other and solve for cos(2x)

OpenStudy (anonymous):

in terms of sin x and cos x

OpenStudy (anonymous):

ok so when i take the derv of sin2x i get 2sinxcosx and the right side is the same ? doesnt make seice

OpenStudy (anonymous):

take the derivative of sin2x using the chain rule

OpenStudy (anonymous):

sin2x = 2sinxcosx; the derivative of sin2x is not equal to 2sinxcosx

OpenStudy (anonymous):

oohh damm yeea its 2cos2x am i right?

OpenStudy (anonymous):

yup :) now take the derivative of 2sinxcosx using the product rule

OpenStudy (anonymous):

would it be -cosxsinx??

OpenStudy (anonymous):

let f(x)=2sin(x) and g(x)=cos(x) the product rule states that the derivative of f(x)g(x) is equal to f(x)g'(x)+f'(x)g(x)

OpenStudy (anonymous):

ok so i get -2sin^(2)x + cos^(2)x

OpenStudy (anonymous):

close. for the second term, remember to carry over the 2 when taking the derivative of 2sin(x)

OpenStudy (anonymous):

like?

OpenStudy (anonymous):

f'(x) * g(x) = [2cos(x)] * [cos(x)] = 2 cos^(2)x

OpenStudy (anonymous):

ohh yeaa

OpenStudy (anonymous):

now what do i do ?

OpenStudy (anonymous):

set the two derivatives equal to each since the functions that we differentiated (sin(x) and 2sin(x)cos(x)) are equal to each other

OpenStudy (anonymous):

and then solve for cos(2x)

OpenStudy (anonymous):

2cos2x = -2sin^(2)x + 2cos^(2)x thus cos(2x) = cos^(2)x-sin^(2)x

OpenStudy (anonymous):

hey here is my new queston. differentitiate the function y=1-cosx/sinx and y=cscx- cotx , and show that the derivaties are equal

OpenStudy (anonymous):

so i toke the derv of the 2ns eqaution and i get the first equation but they say the derv habve to be the same so like what ?

OpenStudy (anonymous):

do i take the derv of them both and they have to end up the same ? ?

OpenStudy (anonymous):

yup, the derivatives should end up the same. also the first function is (1-cosx)/sinx

OpenStudy (anonymous):

can u help me take the derv of the first function , yea ur right

OpenStudy (anonymous):

to take the derivative of (1-cosx)/sinx, use the quotient rule. f(x)=(1-cosx), g(x)=sinx, and the derivative of f/g = (f'(x)g(x)-f(x)g'(x)) / (g(x)^2)

OpenStudy (anonymous):

http://www.1728.org/chainrul.htm

OpenStudy (anonymous):

Ok so this is what i get

OpenStudy (anonymous):

-sin^(2)x-cosx+cox^(2)x/(sinx)^2 right?

OpenStudy (anonymous):

looks good except for the first term. be careful when taking the derivative of 1-cos(x) f'(x) = d/dx (1) + d/dx (-cos(x)) = - d/dx (cos(x)) = sin(x) f'(x)g(x) = sin^(2)x

OpenStudy (anonymous):

ok so then top i get 1-cosx looks good but for my botton its sin^(2)x it should be just sinx thoug

OpenStudy (anonymous):

sin^(2)x is good for the denominator

OpenStudy (anonymous):

but if u take the derv of the 2nd funstion its just sinx for the deno

OpenStudy (anonymous):

if you're allowed to use a trig derivative table d/dx (cscx - cotx) = [-csc(x)cot(x)] - [-csc(x)^2] otherwise, you can use the quotient rule to find the derivatives of csc(x) and cot(x)

OpenStudy (anonymous):

the derv for cscx is 1/sinx and for cotx its cosx/sinx

OpenStudy (anonymous):

cscx = 1/sin(x) and cot(x)=cos(x)/sin(x)

OpenStudy (anonymous):

so it would be 1/sinx-cox/sinx right

OpenStudy (anonymous):

so becasue the deno is the same it would be 1-cosx/sinx right ?

OpenStudy (anonymous):

you need to find the derivatives though. either use a table, or use the quotient rule to find the derivative of 1/sin(x) and the derivative of cos(x)/sin(x)

OpenStudy (anonymous):

dammnn yea

OpenStudy (anonymous):

csc(x)-cot(x)=1/sinx-cox/sinx, since by definition, csc(x)=1/sin(x) and cot(x)=cos(x)/sin(x) here is a trig table: http://sub.allaboutcircuits.com/images/11045.png

OpenStudy (anonymous):

to manually find the derivative of csc(x), which equals 1/sin(x), let f(x)=1 and g(x)=sin(x), then use the quotient rule

OpenStudy (anonymous):

ok man thanks so much

OpenStudy (anonymous):

glad to have helped. let me know if you have any other questions :)

OpenStudy (anonymous):

oo yea i have one more

OpenStudy (anonymous):

here is another one , if f(x)=sinx,evaluate lim h-->0 f(2+h)-f(2)/h , to two decimel places

OpenStudy (anonymous):

^ the question

OpenStudy (isaiah.feynman):